A consumer has an income of $1,000 to spend on food and medicine. The price of one unit of food is $5 and the price of one unit of medicine is $10. For each question, write down the mathematical expression of the budget constraint and draw it carefully. Put food on the x-axis and medicine on the y-axis. Label the incercepts, slopes, and kinks. 1. Suppose the consumer receives coupons for 50 units of food and those can only be used to buy 2. Suppose the consumer receives coupons for 50 units of food and can sell the coupons at half of 3. Instead of the coupon, the consumer receives a 20% discount for additional units of food that food. the market price of the food. exceed 100 units.That is, the consumer pays $5 for each of the first 100 units and $4 for each additional unit.
Answer #1

Answer-Question in the begining provides original budget line i.e., 5x+10y=1000 and further asks how would the equation change and its shiftsMeducu od nal Bu anged p2 hour足し (000- 500 5o0 125 bur much food cah ba pc

Mathematical equations

a. AB 50+10y=1000

BC 5x+10y =1000+50*5

i.e., 5x +10y =1250

b. Shifted Budget line

5x + 10y = 1000+50*2.5 = 1125

c. AB Same as original Budget line.

BD 4x + 10y = 1000